# Audio Array-Based 3D UAV Trajectory Estimation with LiDAR Pseudo-Labeling

### v4

@@ Abstract
Index Terms—  UAV detection, Audio, mel-spectrograms, Self-Supervise, Unsupervised
1 Introduction
In recent years, Unmanned Aerial Vehicles (UAV) [1, 2] have become widely accessible [3], raising significant security concerns such as privacy violations [4, 5], unauthorized surveillance [6, 7, 8], and risks in restricted airspace [9, 10, 11], including threats to civilian aviation. Their use in cross-border drug smuggling [12] and military operations [13, 14] underscores the urgent need for effective UAV tracking technologies [15].
Despite the pressing need for UAV tracking systems, existing studies predominantly rely on visual information [16, 17, 18, 19, 20] for 2D tracking or 3D detection from RADAR/LiDAR [21, 22] [23, 24, 5] [25]. However, these approaches are limited by poor visibility, restricted field of view, and the high cost of RADAR/LiDAR for large-area surveillance and low-altitude smuggling UAV.
     The utilization of audio data, despite its accessibility and cost-effectiveness, remains largely unexplored [26, 27, 28] in UAV trajectory estimation[29, 30]. Given the affordability and ubiquity of microphones compared to LiDAR and cameras [31], there is a significant opportunity to develop innovative 3D tracking solutions [32, 33, 34, 35, 36] that leverage audio information.
@@ 1 Introduction
Fig. 1: System Overview.
@@
Fig. 2: Proposed Self-Supervised Framework.
2 Proposed Method
@@
2.1 Unsupervised Teacher Network
@@
2.2 Self-Supervised Audio Feature Learning
@@
Fig. 3: Estimated Trajectory by our solution
2.3 Gaussian Process Smoothing
@@
Fig. 4: Pham4 Trajectory and Analysis
2.4 Trajectory Estimation Head
@@
Table 1: The comparison of performance. Lower $D_{x,y,z}$ and $E$ represent better trajectory estimation in m.
@@
3 Experiments and Evaluation
3.1 Dataset and Implement Details
@@ 3.2 Evaluation metrics
3.3 Trajectory Estimation Performance
@@
3.4 Benchmark Comparisons
From Table 1, it is clear that our method exhibits a significantly lower average position error compared to all other methods, indicating a superior performance in trajectory prediction. Our methods are compared with two audio-based networks AudioNet and Vora’s Net, two visual-based networks VisualNet and DarkNet, and three audio-visual-based networks ASDNet, AVped, and AV-FDTI.

### v5

@@ Abstract
Index Terms:
UAV detection, Audio, mel-spectrograms, Self-Supervise, Unsupervised
I Introduction
In recent years, Unmanned Aerial Vehicles (UAV) [1, 2, 3] have become widely accessible [4, 5], raising significant security concerns such as privacy violations [6, 7], unauthorized surveillance [8, 9, 10, 11], and risks in restricted airspace [12, 13, 14], including threats to civilian aviation. Their use in cross-border drug smuggling [15] and military operations [16, 17] underscores the urgent need for effective UAV tracking technologies [18].
Despite the pressing need for UAV tracking systems, existing studies predominantly rely on visual information [19, 20, 21, 22, 23, 24, 25] for 2D tracking or 3D detection from RADAR/LiDAR [26, 27] [28, 29, 7, 30, 31, 32, 33, 34, 35]. However, these approaches are limited by poor visibility, restricted field of view, and the high cost of RADAR/LiDAR for large-area surveillance and low-altitude smuggling UAV.
The utilization of audio data, despite its accessibility and cost-effectiveness, remains largely unexplored [36, 37, 38] in UAV trajectory estimation[39, 40]. Given the affordability and ubiquity of microphones compared to LiDAR and cameras [41, 42, 43, 44], there is a significant opportunity to develop innovative 3D tracking solutions [45, 46, 47, 48, 49] that leverage audio information.
@@ I Introduction
Figure 1: System Motivation and Overview.
Figure 2: Proposed Self-Supervised Framework.
@@
II Proposed Method
@@
II-A Unsupervised Teacher Network
@@
II-B Self-Supervised Audio Feature Learning
@@
Figure 3: Estimated Trajectory by our solution
II-C Gaussian Process Smoothing
@@
Figure 4: Pham4 Trajectory and Analysis
II-D Trajectory Estimation Head
@@
TABLE I: The comparison of performance. Lower $D_{x,y,z}$ and $E$ represent better trajectory estimation in m.
@@
III Experiments and Evaluation
III-A Dataset and Implement Details
@@ III-B Evaluation metrics
III-C Trajectory Estimation Performance
@@
III-D Benchmark Comparisons
From Table I, it is clear that our method exhibits a significantly lower average position error compared to all other methods, indicating a superior performance in trajectory prediction. Our methods are compared with two audio-based networks AudioNet and Vora’s Net, two visual-based networks VisualNet and DarkNet, and three audio-visual-based networks ASDNet, AVped, and AV-FDTI.